# The algebraic structure of Dyson–Schwinger equations with multiple insertion places

### v1

@@ Abstract.
KY is supported by an NSERC Discovery grant and by the Canada Research Chairs program. KY thanks Paul Balduf for useful conversations. NOH thanks Lukas Nabergall for sharing code and for many useful discussions in the early stages of this work.
@@ 1. Introduction
The paper proceeds as follows. We will begin by giving background, defining the Connes-Kreimer Hopf algebra and the combinatorial analogues of our Dyson–Schwinger equations in Section 2.1, then defining the Faà di Bruno Hopf algebra and looking in more detail at 1-cocycles in Section 2.2. In Section 2.3 we look at the close relationship between the renormalization group equation and the Riordan group, a relationship which has been underrecognized up to this point. We finally define the Dyson–Schwinger equations in Section 2.4 and give a clean, conceptual way of understanding how the invariant charge relates to the renormalization group equation via our work on the Riordan group. The introduction is then rounded out by Section 2.5 where we consider the anomalous dimension and Section 2.6 where define the tubings that we use for our solutions to Dyson–Schwinger equations and state the previous results in that direction from [tubings].
@@ 2.2. Hopf algebras of polynomials and power series
Let $H$ be a graded connected Hopf algebra and $\phi:H\rightarrow\mathbb{K}[L]$ be an algebra morphism. Then $\operatorname{lin}\phi$ is an infinitesimal character if and only if $\phi|_{L=0}=\epsilon$. Moreover, the following are equivalent:
@@
We know that $\phi$ is a bialgebra morphism if and only if $\phi|_{L=0}=\varepsilon$ and $\frac{d}{dL}\phi=(\operatorname{lin}\phi)*\phi$, where $\operatorname{lin}$ is the map extracting the linear coefficient. Since $\phi$ is an algebra morphism, its behaviour is determined by what it does to the generators, so these are respectively equivalent to $\Phi(x,0)=\varepsilon(\Pi(x))=x$ and
@@
As a special case of this, note that $d\varepsilon\circledast\psi=d\psi$. When $M=H$ we can write $\Lambda\circledast\psi$ using the left action of $H^{*}$ on $H$ described above:
@@ 2.3. The renormalization group equation and the Riordan group
In the quantum theory context, the renormalization group equation 10 is a very important differential equation because it describes how the Green function changes as the energy scale $L$ changes.
@@
The series $\gamma(x)$ is the anomalous dimension. If we make a spacetime dilation $z\mapsto\lambda z$ (with $z$ our space-time point) in a scale invariant quantum field theory then each correlation function changes under the dilation simply by scaling by a constant power of $\lambda$, and that power is known as the scaling dimension. In a non-scale invariant quantum field theory, the correction to such a simple scaling is given by the series $\gamma(x)$. A particularly easy case is when $\beta(x)=0$ in 10 as there the differential equation can be solved by $e^{\gamma(x)L}$ giving a scaling solution at that point.
Returning to our formal series context, taking the coefficient of $L^{0}$ in 10 we see that the anomalous dimension is nothing other than the linear term in $L$ of $G$. Furthermore, writing $G(x,L)=1+\sum_{i\geq 0}g_{i}(x)L^{i}$ (so $\gamma(x)=g_{1}(x)$) and taking the coefficient of $L^{k-1}$ in 10 we obtain
$g_{k}(x)=\frac{1}{k}\left(\beta(x)\frac{d}{dx}-\gamma(x)\right)g_{k-1}(x)$
@@ 2.4. Dyson–Schwinger equations
The diagrammatic form of the Dyson–Schwinger equations mentioned above is perhaps the easiest perspective to get an intuition for what these equations are telling us—they are telling us how to build all Feynman diagrams contributing to a given process by inserting simpler primitive333primitive in a renormalization Hopf algebra Feynman diagrams into themselves. This explains some otherwise mysterious aspects of the nomenclature. The indexing set $I$ for systems is typically giving the external edges of the diagram. The weight is the loop order (dimension of the cycle space) of the primitive diagram. The insertion exponent counts how the number of places a Feynman diagram of the given external edge structure can be inserted grows as the loop order grows. The function $A(\rho)/\rho$ is the Mellin transform of the Feynman integral for the primitive diagram regularized at the insertion place.
@@ 2.6. Tubings
We now introduce the combinatorial objects that we will use to understand 1-cocycles. We will only be interested in trees, but the basic definitions can be given in the context of an arbitrary finite poset $P$. A tube is a connected convex subset of $P$ [^5] Note that being both connected and convex is equivalent to inducing a connected subgraph of the Hasse diagram of $P$. For tubes $X,Y$ write $X\to Y$ if $X\cap Y=\emptyset$ and there exist $x\in X$ and $y\in Y$ such that $x<y$. A collection $\tau$ of tubes is called a tubing if it satisfies the following conditions:
@@ 3.1. Context
Finally, to each $p$ we associate a vector of indeterminates $\mathbf{L}_{p}=(L_{e})_{e\in E_{p}}$ and a 1-cocycle $\Lambda_{p}\in\mathrm{Z}^{1}(\mathbb{K}[\mathbf{L}_{p}],\mathbb{K}[L])$. The Dyson–Schwinger associated to these data is
@@ 4. Conclusion
Paul Balduf has done some numerical computations of solutions to Dyson–Schwinger equations with two insertion places (see pp. 44, 45 of [Bslides]). Interestingly, when the two insertion places degenerate into one by setting the power of one of the insertions (the parameter $s$) to $-1$, then the function does not appear to be differentiable based on the result of the numerical computation. It is not clear why this should be, and we certainly have no combinatorial understanding in this direction at present.

### v2

@@ Abstract.
KY is supported by an NSERC Discovery grant and by the Canada Research Chairs program. KY thanks Paul Balduf for useful conversations. NOH thanks Lukas Nabergall for sharing code and for many useful discussions in the early stages of this work. Thanks also to the referee for their suggestions.
@@ 1. Introduction
The paper proceeds as follows. We will begin by giving background, defining the Connes-Kreimer Hopf algebra and the combinatorial analogues of our Dyson–Schwinger equations in Section 2.1, then defining the Faà di Bruno Hopf algebra and looking in more detail at 1-cocycles in Section 2.2. In Section 2.3 we look at the close relationship between the renormalization group equation and the Riordan group, a relationship which has been underrecognized up to this point. We finally define the Dyson–Schwinger equations in Section 2.4 and give a clean, conceptual way of understanding how the invariant charge relates to the renormalization group equation via our work on the Riordan group. The introduction is then rounded out by Section 2.5 where we consider the anomalous dimension and Section 2.6 where we define the tubings that we use for our solutions to Dyson–Schwinger equations and state the previous results in that direction from [tubings].
@@ 2.2. Hopf algebras of polynomials and power series
Let $H$ be a graded connected Hopf algebra and $\phi:H\rightarrow\mathbb{K}[L]$ be an algebra morphism. Then $\operatorname{lin}\phi$ is an infinitesimal character if and only if $\phi|_{L=0}=\epsilon$, where $\operatorname{lin}$ is the map extracting the linear coefficient. Moreover, the following are equivalent:
@@
Recall the notation and results of Item 2.3(iii). We know that $\phi$ is a bialgebra morphism if and only if $\phi|_{L=0}=\varepsilon$ and $\frac{d}{dL}\phi=(\operatorname{lin}\phi)*\phi$. Since $\phi$ is an algebra morphism, its behaviour is determined by what it does to the generators, so these are respectively equivalent to $\Phi(x,0)=\varepsilon(\Pi(x))=x$ and
@@
{example}
For example, take $\phi\colon\mathsf{FdB}\to\mathbb{K}[L]$ defined on generators by $\phi(\pi_{n})=L^{n}$ and extended as an algebra morphism. Then $\Phi(x,L)=\sum_{n=0}^{\infty}L^{n}x^{n+1}=\frac{x}{1-Lx}$ and $\beta(x)=x^{2}$. We can now check the conditions from Theorem 2.4: we have $\Phi(x,0)=x$ and $\frac{\partial\Phi(x,L)}{\partial L}=x^{2}\sum_{n=1}^{\infty}nL^{n-1}x^{n-1}=\beta(x)\frac{\partial\Phi(x,L)}{\partial x}$, hence $\phi$ is a bialgebra morphism.
@@
As a special case of this, note that $d\varepsilon\circledast\psi=d\psi$ (dropping the subscript to the coboundary map by the usual abuse of notation). When $M=H$ we can write $\Lambda\circledast\psi$ using the left action of $H^{*}$ on $H$ described above:
@@ 2.3. The renormalization group equation and the Riordan group
In the quantum field theory context, the renormalization group equation 10 is a very important differential equation because it describes how the Green function changes as the energy scale $L$ changes.
@@
The series $\gamma(x)$ is the anomalous dimension. It is a dimension in the sense of scaling dimension and it is anomalous in the sense that in the classical setting one would have a constant integer in place of $\gamma(x)$. A particularly easy case is when $\beta(x)=0$ in 10 as there the differential equation can be solved by $e^{\gamma(x)L}$ which is of a particularly simple form sometimes called a scaling solution.
Returning to our formal series context, taking the coefficient of $L^{0}$ in 10 we see that the anomalous dimension is nothing other than the linear term in $L$ of $G$ (in the cases of physical interest we will always have $g_{0}=1$). Furthermore, writing $G(x,L)=1+\sum_{i\geq 0}g_{i}(x)L^{i}$ (so $\gamma(x)=g_{1}(x)$) and taking the coefficient of $L^{k-1}$ in 10 we obtain
$g_{k}(x)=\frac{1}{k}\left(\beta(x)\frac{d}{dx}+\gamma(x)\right)g_{k-1}(x)$
@@ 2.4. Dyson–Schwinger equations
The diagrammatic form of the Dyson–Schwinger equations mentioned above is perhaps the easiest perspective to get an intuition for what these equations are telling us—they are telling us how to build all Feynman diagrams contributing to a given process by inserting simpler primitive333primitive in a renormalization Hopf algebra, or from a physics perspective, subdivergence-free Feynman diagrams into themselves. This explains some otherwise mysterious aspects of the nomenclature. The indexing set $I$ for systems is typically giving the external edges of the diagram. The weight is the loop order (dimension of the cycle space) of the primitive diagram. The insertion exponent counts how the number of places a Feynman diagram of the given external edge structure can be inserted grows as the loop order grows. The function $A(\rho)/\rho$ is the Mellin transform of the Feynman integral for the primitive diagram regularized at the insertion place.
@@ 2.6. Tubings
We now introduce the combinatorial objects that we will use to understand 1-cocycles. We will only be interested in trees, but the basic definitions can be given in the context of an arbitrary finite poset $P$. A tube is a connected convex subset of $P$. For tubes $X,Y$ write $X\to Y$ if $X\cap Y=\emptyset$ and there exist $x\in X$ and $y\in Y$ such that $x<y$. A collection $\tau$ of tubes is called a tubing if it satisfies the following conditions:
@@
The linearized coproduct is itself a nice algebraic object as it is co-pre-Lie and specifically is dual to the pre-Lie product given by insertion of rooted trees. See [CLpre] for more on the structure of this product.
@@ 3.1. Context
Finally, to each $p$ we associate a vector of indeterminates $\mathbf{L}_{p}=(L_{e})_{e\in E_{p}}$ and a 1-cocycle $\Lambda_{p}\in\mathrm{Z}^{1}(\mathbb{K}[L],\mathbb{K}[\mathbf{L}_{p}])$. The Dyson–Schwinger associated to these data is
@@
Very similar combinatorial and algebraic structures are used in the study of regularity structures, see [BHZregularity].
@@ 4. Conclusion
Paul Balduf has done some numerical computations of solutions to Dyson–Schwinger equations with two insertion places. See pp. 44, 45 of [Bslides] where he plots the two parameters controlling the exponential growth rate of the coefficients of the anomalous dimension (what he calls $\mu$ and $\lambda$ in the form $(-\lambda)^{-n}\Gamma(n-\mu)$) as a function of the power of the insertions (in our notation $\mu_{1}-1$ and $\mu_{2}-1$). Interestingly, when the two insertion places degenerate into one by setting the power of one of the insertions to $-1$ (that is, one of the $\mu_{i}=0$), then the function does not appear to be differentiable based on the result of the numerical computation. It is not clear why this should be, and we certainly have no combinatorial understanding in this direction at present.
@@
The way in which trees with edge decorations and their Hopf algebraic structures appear in the study of regularity structures, as described in [BHZregularity], has a very similar flavour to what we see in our work on Dyson–Schwinger equations. It would be very interesting to investigate a more precise connection between these areas.